# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Appendix D Observable and Correlation Functions
where $c_{i\uparrow}^{\dagger}$ creates an up-spin fermion on site $i$ and $c_{i\downarrow}^{\dagger}$ creates a down-spin fermion on site $i$. The spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ can be written in terms of these fermion operators as:
@@
SII Staggered Heisenberg Order Parameter
Here we express the staggered-Heisenberg order parameter in the hybrid $(x,k_{y})$ operator representation, where $x$ indexes cylinder rings. Because of translation invariance around the cylinder, it is useful to define the $y$-averaged inter-ring NN Heisenberg operator
@@
Since both the IQH and CSL ground states exhibit a two-ring translation invariance (i.e., are invariant under $(T_{x})^{2}$ in addition to $T_{y}$) then we can measure the local order parameter
@@
$=\frac{1}{L_{y}^{2}}\sum_{a=x,y,z}\sum_{k,k^{\prime}}\sum_{p,p^{\prime}}\left(\frac{1}{L_{y}}\sum_{y}e^{i(k-k^{\prime}+p-p^{\prime})y}\right)c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime})\cdot c^{\dagger}(x+1,p)\frac{\sigma^{a}}{2}c(x+1,p^{\prime})$
@@
Appendix E Correlation Lengths
Figure S3: Correlation length data for even-circumference geometries. Left panel: Illustration of the external magnetic flux penetrating each loop of the infinite cylinder, namely no flux, $\Phi=0$. Center panels: Maximum correlation length at $L_{y}=4$ in the $Q=0e$ charge sector [blue] for a range of bond dimensions increasing with increasing darkness; and in the $Q=1e$ charge sector [green]. Right panel: The same for $L_{y}=6$.
Figure S4: Correlation length data for even-circumference geometries. Left panel: Illustration of the external magnetic flux penetrating each loop of the infinite cylinder, which alternates between ±$\pm\pi/2$, corresponding to threaded flux $\Phi_{\text{ext}}=\pi/2$ in our conventions. Center panels: Maximum correlation length at $L_{y}=3$ in the $Q=0e$ charge sector [blue] for a range of bond dimensions increasing with increasing darkness; and in the $Q=1e$ charge sector [green]. Right panel: The same for $L_{y}=5$.
@@ Appendix E Correlation Lengths
Here we briefly comment on additional data concerning the correlation lengths for geometries not displayed in the text. In particular, Fig. S3 shows correlation length data for the $L_{y}=4$ and $6$ cylinders which are threaded by no external flux. In particular, $L_{y}=6$ reproduces the behavior presented in Kuhlenkamp et al., [11]. We remark that the correlation lengths are quantitatively smaller than those at $\Phi_{\text{ext}}$ external flux, which are shown in the main text. We similarly show data for the $L_{y}=3$ and $5$ cylinders threaded by $\Phi_{\text{ext}}=\pi/2$ external flux. Unlike at $\Phi_{\text{ext}}=0$ in the main text, the flux breaks the glide-PH symmetry, and thus we see no diverging or quantitatively large correlation length (as is consistent with no spontaneous symmetry breaking).
Appendix F Correlation Length Extrapolation
@@
Figure S5: (a) Extrapolated inverse correlation length in the charge-neutral sector at $L_{y}=4$ and $\Phi_{\text{ext}}=\pi/2$ obtained following the method in Ref. [52], as detailed in Sec. F, for a range of Hubbard strengths [different colors] and for a sequence of increasing bond dimensions [same-colored points moving sequentially to the left]. The extrapolated inverse correlation length saturates at a minimum value exceeding $0.15$; the correlation length is equivalently bounded above by $6.6$ ring spacings. (b) The same for $L_{y}=5$ and $\Phi_{\text{ext}}=0$. In this case, the extrapolated inverse correlation length reaches and descends beyond 0, signaling the need for larger bond dimensions to accurately estimate large correlation lengths.
@@
To obtain the extrapolated correlation lengths shown in the main text, we employ a recently-developed method for extrapolating correlation lengths in DMRG [52, 59]. Namely, we compute several of the largest eigenvalues of the transfer matrix $TX_{i}=\lambda_{i}X_{i}$ in the electric charge $Q_{E}=1$ sector for each $k_{y}$, and select those from the same “excitation branch” [52], namely carrying a consistent complex phase. These are related to the correlation lengths as
$\epsilon_{i}=1/\xi_{i}=-\log\left|\lambda_{i}\right|$ (82)
@@
with $\epsilon_{i}\leq\epsilon_{i+1}$. We report $\xi$ in units of cylinder rings, or equivalently $u=a\sqrt{3}/2$ where $a$ is the lattice spacing of the underlying triangular lattice. Correspondingly, $\epsilon$ is reported in units of $u^{-1}$. For a system that is gapped in a charge sector $q=(Q,S_{z},k_{y})$, both $\epsilon_{1}^{q},\epsilon_{2}^{q}$ converge to the same infinite value as $\chi\to\infty$, namely the finite inverse correlation length. As a result, the quantity
$\delta^{q}=\epsilon_{2}^{q}-\epsilon_{1}^{q}$ (83)
@@
vanishes at $\chi\to\infty$, serving as a good “scaling variable” that measures the deviation from convergence. To estimate $\xi(\chi\to\infty)$, we therefore compute $\epsilon(\delta)$ at each available $\epsilon(\chi),\delta(\chi)$ and extrapolate to $\delta=0$ with a polynomial fit, namely
@@
This extrapolation is shown in Fig. S5 for both $L_{y}=4$ at $\Phi_{\mathrm{ext}}=\pi/2$ and $L_{y}=5$ at $\Phi_{\mathrm{ext}}=0$, yielding the extrapolated values shown in the main text. In both cases, the correlation length is computed in the $(Q,S_{z},k_{y})=\bm{0}$, which hosts largest correlation length.
Appendix G Density-density correlations at all circumferences
Figure S6: Density-density connected correlations for each of the four studied circumferences $L_{y}=3$ to $6$, with the Hubbard interaction tuned to the values at which the correlation length is maximal at the largest available bond dimension, namely $U(L_{y}=3)=11.125$, $U(L_{y}=4)=11.0$, $U(L_{y}=5)=11.6$, and $U(L_{y}=6)=11.4$, respectively. The even-circumference cylinders are threaded with $\Phi_{\text{ext}}=\pi/2$, as in the main text.
In Fig. S6 we show the ring-averaged density-density connected correlations, defined as $\langle n_{\text{av}}(x)n_{\text{av}}(0)\rangle,$ where
$n_{\text{av}}(x)$ $=\frac{1}{L_{y}}\sum_{y}c^{\dagger}(x,y)c(x,y)=\frac{1}{L_{y}}\sum_{k}c^{\dagger}(x,k)c(x,k).$ (85)
Panels (a,c), which correspond to the odd circumferences $L_{y}=3,5$ respectively, exhibit apparently algebraic correlations at long distances. On the other hand, panels (b,d) exhibit exponentially-decaying correlations that nonetheless feature a near-algebraic plateau, over a range of $x$ less than their respective correlation lengths.

### v2

@@ Appendix E Operators in the hybrid $x\,k_{y}$ representation
The spin correlation operator $\vec{S}_{i}\cdot\vec{S}_{j}$ can then be written in terms of these fermion operators as:
@@
Next, we express the staggered-Heisenberg order parameter in the hybrid $(x,k_{y})$ operator representation, where $x$ indexes cylinder rings. Because of translation invariance around the cylinder, it is useful to define the $y$-averaged inter-ring NN Heisenberg operator
@@
$=\frac{1}{L_{y}^{2}}\sum_{a=x,y,z}\sum_{k,k^{\prime}}\sum_{p,p^{\prime}}\left(\frac{1}{L_{y}}\sum_{y}e^{i(k-k^{\prime}+p-p^{\prime})y}\right)$
$\qquad\quad\times c^{\dagger}(x,k)\frac{\sigma^{a}}{2}c(x,k^{\prime})\cdot c^{\dagger}(x+1,p)\frac{\sigma^{a}}{2}c(x+1,p^{\prime})$
@@
Figure S5: Left panel: Sketch of three rings of cylinder at indicated $x$ values (cylinder ring indices), with solid and dashed lines showing along which staggered-Heisenberg order parameter [defined in main text] changes under glide particle-hole symmetry $PT_{x}$ on odd-circumference cylinders. Right panel: Magnitude of the staggered Heisenberg order parameter of Eq. (84) plotted against $U/t$ for cylinder widths indicated and with $PT_{x}$ preserving and broken phases indicated by the check and cross symbols. Consistent with the recovery of this symmetry at the putative 2+1D critical point, the magnitude of the order parameter [bottom left] decreases substantially from $L_{y}=3$ to 5.
Since both the IQH and CSL ground states exhibit a two-ring translation invariance (i.e., are invariant under $(T_{x})^{2}$ in addition to $T_{y}$) then we can measure the following operator defined on the first two bonds near the origin:
In Fig. S5, we plot this Heisenberg order parameter on the $L_{y}=3,5$ geometries with $\Phi_{\text{ext}}=0$, which have the glide-particle-hole symmetry. This is an alternative to the order parameter computed and displayed in in Fig. 3(a) of the main text, and defined in Eq. (3), which is instead based on the circumferential current operator. Both order parameters demonstrate that $\mathcal{P}T_{x}$ symmetry is spontaneously broken in the CSL phase on these two geometries.
Appendix F Even-circumference supplementary data
SI Reproducing correlation lengths of Kuhlenkamp et al.
Figure S6: (a) Plot of the maximum correlation length in the charge-$0e$ [middle panel] and charge-$1e$ [right panel] sectors of the ground state transfer matrix for the YC6 geometry (at the particle-hole-symmetric value $\Phi_{\text{ext}}=0$), with bond dimensions increasing light-to-dark [see legend]. Left illustration: depiction of the flux through the various cylinder rings, consistent with no externally-threaded magnetic flux. (b) Same for YC4, likewise at $\Phi_{\text{ext}}=0$.
@@ Appendix F Even-circumference supplementary data
Here we briefly comment on additional data concerning the correlation lengths for geometries not already highlighted in the main text. In particular, Fig. S6 shows correlation length data for the $L_{y}=4$ and $6$ cylinders which are not threaded by external flux.
In particular, $L_{y}=6$ reproduces the behavior presented in Kuhlenkamp et al., [19], pushed here to bond dimension $\chi=32768$. We remark that the correlation lengths are smaller than those with $\Phi_{\text{ext}}=\pi/2$ external flux exhibited in the main text, but similarly exhibit a maximum in the correlation length in the $(Q,S_{z})=(0e,0\hbar)$ sector. The peak (whose height is still increasing with bond dimension) is located at $U/t=10.6$ at the largest bond dimension $\chi=32768$; the $U$ location of the peak is also still slowly growing with bond dimension.
On the $L_{y}=4$ cylinder, the correlation length in the $(Q,S_{z})=(0e,0\hbar)$ sector similarly exhibits a finite, local maximum. It is peaked at the smaller value $U/t=10.0$. There is an uptick in the correlation length at $U/t=11.7$, which is accompanied by an identical value in the $(Q,S_{z})=(0e,\pm\hbar)$ sectors; this feature indicates the softness of a spin-triplet excitation at $L_{y}=4$, and is sub-dominant in comparison at $L_{y}=6$. We note that the $Q=1e$ correlation length does not show any feature in the vicinity of this peak or at the uptick.
SII Flux-threaded system: correlation length extrapolation for main text
@@
Figure S7: (a) Inverse correlation length in the charge-neutral sector at $L_{y}=4$ and $\Phi_{\text{ext}}=\pi/2$. The $y$-intercept of the fitted curve is the correlation length extrapolated to infinite bond dimension; the curve is obtained following the method in Ref. [64], as detailed in Sec. SII, for a range of Hubbard strengths [different colors], sampled from a sequence of increasing bond dimensions [same-colored points moving sequentially to the left]. The extrapolated inverse correlation length saturates at a minimum value exceeding $0.15$, i.e., the correlation length is bounded above by $6.6$ ring spacings. (b) The same for $L_{y}=5$ and $\Phi_{\text{ext}}=0$.
@@
To obtain the extrapolated correlation lengths shown in the main text, we employ a recently-developed method for extrapolating correlation lengths in DMRG [64, 82]. Namely, we compute several of the largest eigenvalues of the transfer matrix $TX_{i}=\lambda_{i}X_{i}$ in a given charge (i.e., electric charge and spin) sector for each $k_{y}$, and select those from the same “excitation branch” [64], namely carrying a consistent complex phase. These are related to the correlation lengths by
$\epsilon_{i}=1/\xi_{i}=-\log\left|\lambda_{i}\right|,$ (85)
@@
with $\epsilon_{i}\leq\epsilon_{i+1}$. Throughout this work, we report $\xi$ in units of cylinder rings, or equivalently $u=a\sqrt{3}/2$ where $a$ is the lattice spacing of the underlying triangular lattice. Correspondingly, $\epsilon$ is reported in units of $u^{-1}$. For a system that is gapped in a charge sector $q=(Q,S_{z},k_{y})$, both $\epsilon_{1,q},\epsilon_{2,q}$ converge to the same finite value as $\chi\to\infty$, namely the finite inverse correlation length. As a result, the quantity
$\delta_{q}=\epsilon_{2,q}-\epsilon_{1,q}$ (86)
@@
vanishes at $\chi\to\infty$, serving as a good “scaling variable” that measures the deviation from convergence. To estimate $\xi(\chi\to\infty)$, we therefore compute $\epsilon(\delta)$ at each $\chi$ (which gives a data point $(\epsilon(\chi),\delta(\chi))$) and extrapolate to $\delta=0$ by fitting to the form
@@
This extrapolation is shown in Fig. S7 for $L_{y}=4$ at $\Phi_{\mathrm{ext}}=\pi/2$, yielding the extrapolated values shown in the main text. The correlation length is computed in the $(Q,S_{z},k_{y})=0$ charge sector, which hosts largest correlation length. We also perform the extrapolation for $L_{y}=5$ at the glide-particle-hole symmetric flux $\Phi_{\mathrm{ext}}=0$, which instead exhibits a divergence in $\xi_{0}$ at a value of $U$ in the range $U\in(11.4,11.8)$; In the main text we take $U_{c}(L_{y}=6)=11.6$, at which the correlation length is largest at the largest bond dimension $\chi=16384$.
Appendix G Supplementary odd-circumference analysis
SI YC3 Correlation Lengths
Figure S8: (a) Plot of the maximum correlation length in the charge-$0e$, charge-$1e$, and charge-$2e$ sectors of the ground state transfer matrix for the YC3 cylinder geometry, with bond dimensions increasing light-to-dark. Left illustration: depiction of the flux through the various cylinder rings, consistent with no externally-threaded magnetic flux, $\Phi_{\text{ext}}=0$. (b) Same for the case $\Phi_{\text{ext}}=\pi/2$, for which the Hamiltonian explicitly breaks glide-PH symmetry and therefore does not exhibit SSB criticality.
In Fig. S8, we display the correlation lengths for the YC3 geometry. In particular, in panel (a) we provide data for the $\Phi_{\text{ext}}=0$ case, which accompanies Fig. 3 of the main text. Here, the system respects glide-PH symmetry and, like the YC5 system with $\Phi_{\text{ext}}=0$ studied in the main text, exhibits a symmetry-unbroken small-$U$ phase and symmetry-broken larger-$U$ phase. The transition between then is evidently continuous, featuring a diverging correlation length which reaches almost 600 ring spacings at the maximum bond dimension $\chi=16384$ considered. In Fig. S8(b), in contrast, threading $\Phi_{\text{ext}}$ explicitly breaks the glide-PH symmetry, which reduces the transition to a crossover with a small and wide peak in the correlation length.
SII Relation to ionic Hubbard chain
Here, we make more precise the relationship between the “ionic Hubbard” chain and our model on those YC-$L_{y}$ cylinders (with odd $L_{y}$ and $\Phi_{\text{ext}}=0$) which possess glide-PH symmetry. In particular, the former model can be thought of as a truly 1+1D version of the latter, i.e., a bipartite chain as opposed to a quasi-(1+1)D system with the hopping-connectivity of a cylinder. We first compare them at the level of symmetries. The ionic Hubbard Hamiltonian is defined as [83]:
$H=-t\sum_{i,\sigma}(c^{\dagger}_{i,\sigma}c_{i+1,\sigma}+h.c.)+U\sum_{i}n_{i,\uparrow}n_{i,\downarrow}+\frac{\Delta}{2}\sum_{i,\sigma}(-1)^{i}n_{i,\sigma}.$ (88)
The Hamiltonian is invariant under translation by two sites, with $\Delta$ being the potential energy difference between the two sublattice sites in the unit cell. Moreover, it is invariant under site-centered inversion (but not bond-centered inversion), time-reversal, and a “duality” symmetry that is the combination of particle-hole with a single-site translation [73]. These symmetries of the ionic Hubbard chain correspond to the following symmetries of our model: $(T_{x})^{2},$ site-centered $C_{2z}$, $M_{y}\mathcal{T}$, and glide-PH $\mathcal{P}T_{x}$.
Remarkably, the two models are also similar at the level of their ground states. The small-$U$ ground state phase of the ionic Hubbard chain is a fully-symmetric band insulator. Is it widely believed that the phase which appears next upon increasing $U$ is a “spontaneously-dimerized” phase which spontaneously breaks the duality and site-centered inversion symmetries [73]. Moreover, as mentioned in the main text, it is believed that these two states are separated by a 2D Ising critical point. This Ising picture has been substantiated both by weak-coupling bosonization arguments [70, 71] and an effective strong-coupling effective spin model formulation [72]. However, at least one careful numerical study of the transition using finite DMRG finds features inconsistent with Ising criticality [84], though the numerical observations are still consistent with a gap to charged excitations and a gapless neutral sector [85]. These two phases should be identified with the odd-circumference (symmetry-respecting) IQH and (symmetry-breaking) CSL phases, respectively.
Appendix H Density-density correlations at all circumferences
In Fig. S9 we show various connected correlation functions for the even-circumference cylinders studied in Fig. 3 of the main text, namely $L_{y}=4,6$ with $\Phi_{\text{ext}}=\pi/2$. We plot the current-current (Fig. S9a) and density-density (Fig. S9b) connected correlation functions, which extend far beyond the spin-spin (Fig. S9c) and electron-hole (Fig. S9d) correlators (defined in main text). The first two extend to much longer range than the latter, suggesting both the existence of a spin gap across the transition and the prominent of spin-singlet charge-neutral fluctuations at criticality. While the even-$L_{y}$ systems always exhibit a crossover from IQH to CSL, their behavior is qualitatively consistent with the odd-$L_{y}$ cylinders with $\mathcal{P}T_{x}$ symmetry: the spin-spin and electron-hole correlations are also rapidly exponentially decaying, and the current-current correlator is likewise dominant.
Figure S9: Connected correlation functions for the $L_{y}=4,6$ systems with $\Phi_{\text{ext}}=\pi/2$. The Hubbard interaction is tuned to the value where the correlation lengths are maximal, namely $U(L_{y}=4)=11.0$ and $U(L_{y}=6)=11.4$. (a) Current-current correlator (defined in main text). (b) Correlator of ring-averaged charge density. (c) Correlator of ring-averaged spin density density. (d) Electron-hole correlator (defined in main text). The $L_{y}=6$ correlations are reduced by $\times 10$ for ease of viewing.